OpenStudy (anonymous):

Write the following in standard form using integers. y = 2/3x -7

OpenStudy (anonymous):

the standard equation is Ax+By=C \[y=\frac{ 2 }{ 3 }x-7\] \[-\frac{ 2 }{ 3 } x+y=-7\] then multiply both sides by -1 so... \[\frac{ 2 }{ 3 }-y=7\]